4 persons are choosen at random from a group of 3 men, 2 women, 3 children. Find the probability
that out of 4 choice, exactly 2 are children.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Probability that out of 4 exactly 2 children is = (3c2)(5c2)/8c4
=30/70=3/7
Thanks and Regards
Shaik Aasif
askIITians faculty